OpenStudy (anonymous):

help plse: (4a-3) = a+33/a+1 solution is

OpenStudy (anonymous):

help plse: (4a-3) = a+33/a+1 solution is

OpenStudy (zarkon):

is the equation \[4a-3=\frac{a+33}{a+1}\]?

OpenStudy (anonymous):

yes where 4a-3 in parenthese

OpenStudy (zarkon):

multiply by a+1 \[(4a-3)(a+1)=a+33\] multiply out \[4a^2+4a-3a-3=a+33\] \[4a^2+a-3=a+33\] \[4a^2-3=33\] \[4a^2=36\] \[a^2=9\] \[a=\pm3\]

OpenStudy (anonymous):

so the answer would be 3

OpenStudy (zarkon):

3 and -3

OpenStudy (anonymous):

Zarkon can ou help with this one 2-a-4/a+3 = a^2-2/a+3

OpenStudy (zarkon):

you need to work on formatting your questions. The way you type them is very ambiguous. is this your equation either of these? \[2-a-\frac{4}{a+3}=\frac{a^2-2}{a+3}\] \[2-\frac{a-4}{a+3}=\frac{a^2-2}{a+3}\]

OpenStudy (anonymous):

it is the 2nd one I'm sorry for the formatting I do not know how to use the equation format too goo so I did the best I could sorry for the confusion.l

OpenStudy (zarkon):

ok \[2-\frac{a-4}{a+3}=\frac{a^2-2}{a+3}\] \[2=\frac{a^2-2}{a+3}+\frac{a-4}{a+3}\] \[2=\frac{a^2-2+a-4}{a+3}\] \[2(a+3)=a^2-2+a-4\] \[2a+6=a^2-2+a-4\] \[0=a^2-2+a-4-2a-6\] \[0=a^2-a-12\] \[0=(a-4)(a+3)\] so a=4 -3 is not a solution since it does not satisfy the original equation (can't divide by 0)